OpenStudy (anonymous):

A motorboat takes 5 hours to travel 400mi going upstream. The return trip takes 2 hours going downstream. What is the rate of the boat in still water and what is the rate of the current?

OpenStudy (anonymous):

I set up trip 1 being 400=(B-C)5h and trip 2 being 400=(B+C)2h but I keep coming up with weird numbers so I must be doing something wrong

OpenStudy (anonymous):

set the distance for both trips equal. 5(velocity of boat in still water-current)=2(velocity of boat+current) 3velocity=7current velocity=(7/3)current now plug into either 2(7/3 +3/3)current=400 20/3(current)=400 current/3=20 current=60 The rest is easy.

OpenStudy (anonymous):

Ok.. I think I got it. I had velocity = (7/5) current and was screwed from there. Thank you for your help!

OpenStudy (anonymous):

I came up with speed of the boat would be 140 and your answer of current is 60.. would that be correct?